OpenStudy (ssddss02):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay? A. B. C. D.

OpenStudy (anonymous):

120(9/100)=?

OpenStudy (ssddss02):

A. 9/120 = x/100 B. 120/x = 100/9 C. x/120 = 100/9 D. 120/9 = 100/x Those are the choices

OpenStudy (ssddss02):

Help!

OpenStudy (texaschic101):

i am thinking b

OpenStudy (ssddss02):

I think its B.

OpenStudy (anonymous):